# Leveraging priors on distribution functions for multi-arm bandits

### v1

@@ 6.1 Admissible probability matching algorithms
It’s easy to notice in the preceding analysis that there’s no restriction on $\mathbb{P}_{t}(A^{\star}=a)$ to be derived using a Bayes-rule based posterior-distributions of arm-rewards,$\mathbb{P}_{t}(R_{t,a})$ as is done in parametric Thompson sampling. This choice is rather implicit, given the decision theoretic and information theoretic coherency of Bayesian framework (Wald, 1961; Zellner, 1988). However, Bayesian-framework is not limited to Bayes-rule based derivation of posterior distributions. Another valid Bayesian approach (Orbanz, 2009; Ghosal & van der Vaart, 2017) for obtaining posteriors is leveraging the property of conjugacy as discussed in Sec 3. In particular, most nonparametric priors do not satisfy the necessary conditions for Bayes rule (See  A.1), and one must rely on their conjugacy property to derive the corresponding posteriors. Therefore, all probability matching algorithms which derive $\mathbb{P}_{t}(R_{t,a})$ (and hence $\mathbb{P}_{t}(A^{*}=a)$) using a valid Bayesian approach are admissible in the information theoretic analysis of Russo & Van Roy (2014). Additionally, these admissible algorithms would enjoy similar bounds as parametric Thompson sampling on their information-ratio (and consequently Bayesian regret), if they satisfy auxiliary conditions required from the original analysis.
For the case of $\sigma$-sub Gaussian reward noise discussed before, it is easy to see that we require the following auxiliary conditions: In each round $t$, (1) the instantaneous reward noise, $X(t)$, in Lemma 5, is $\sigma$-sub-Gaussian; (2) $\texttt{KL}(\mathbb{P}_{t}(R_{t,a}|A^{\star}=a)||\mathbb{P}_{t}(R_{t,a}))$ in Lemma 5 is well defined. The second condition holds if $\mathbb{P}_{t}(A^{\star}=a)>0$ owing to a classical fact in conditional probability (Williams, 1991),
Fact 6.
@@
For any random variable $Z$ and event $E\subset\Omega$, where $\Omega$ is the probability space, if $\mathbb{P}_{t}(E)=0$, then $\mathbb{P}_{t}(E|Z)=0$ almost surely. Conversely, for any $x\in\mathcal{X}$ with $\mathbb{P}_{t}(X=x)>0$, $\mathbb{P}_{t}(Y|X=x)$ is absolutely continuous with respect to $\mathbb{P}_{t}(Y)$.
@@
DPPS satisfies all the conditions above: It is admissible since it utilizes a valid Bayesian approach, i.e. conjugacy of DP priors/posteriors, to derive $\mathbb{P}_{t}(A^{\star}=a)$; Also, clearly, $\mathbb{P}_{t}(A^{\star}=a)>0$ whenever the base measure, $F_{0}$, of the DP prior (and hence of the corresponding DP posterior), $\texttt{DP}(\alpha,F_{0})$, is non-null. Finally, the following property of the tail of DP priors/posteriors ensures $\sigma$-sub-Gaussian nature of the instantaneous reward noise, $X(t)$, whenever the base measure, $F_{0}$, of the DP prior, $\texttt{DP}(\alpha,F_{0})$, is $\sigma$-sub-Gaussian,
Fact 7 (Doss & Sellke (1982)).
@@
This leads us to the following upper bound on Bayesian regret of DPPS,
@@
For the setting of $\sigma$-sub-Gaussian rewards, starting with a DP-prior with a $\sigma$ sub-Gaussian base measure, the Bayesian regret of DDPS satisfies
@@ 7 Conclusions and Perspectives
Firstly, we point that DPs are not the only Bayesian nonparametric priors on the space of distribution functions, and further generalization of DPPS is possible. For example, other probability matching algorithms using Pitman-Yor (Pitman & Yor, 2006) processes and Pólya-Tree priors  (Castillo, 2017; 2024) can be useful generalizations of DPPS. Next note that, although we derived DPPS for multi-arm bandits without any structure, we believe the results in this paper would carry out on other types of online learning problems studied in  (Russo & Van Roy, 2016), e.g. linear bandits. Also, since all the Bayesian regret guarantees of Thompson sampling in (Russo & Van Roy, 2016) hold for Information directed sampling (IDS) (Russo & Van Roy, 2014), we conjecture that a DPPS version of IDS would also be optimal following the arguments in our paper. This can be useful since IDS has been specifically shown to be asymptotically optimal for problems wherein Thompson sampling and UCB type algorithms fail  (Lattimore & Szepesvari, 2017) to be so. A major hurdle in IDS is however its computational-complexity, owed to intractable posteriors that result because of the use of parametric-posteriors based on Bayes-rule. It would be interesting, in future work, to study a nonparametric variant of IDS that utilizes DP posteriors as it would overcome these computational issues.
Finally, we consider DPPS as a generic design principle, based on Bayesian non-parametric statistics, that can be extended to the setting of Markov Decision Processes (MDPs) as well. This can be done in both model-based and model-free scenarios. In the former, a Posterior Sampling Reinforcement Learning (PSRL) (Osband et al., 2013; Fan & Ming, 2021) algorithm based on Dirirchlet Process posteriors is definitely a promising direction of research. For the model-free scenario, one can extend Randomized Least Square Value Iteration (RLSVI) from its current Bayesian-Bootstrap based implementations (Osband et al., 2016; 2019) to a full-fledged DP implementation to account for uncertainty that does not come from the observed data, in a principled manner similar to that shown in this paper. We leave these intriguing research questions and extensions for future work.
@@ Acknowledgments
This work has been supported by the French Ministry of Higher Education and Research, the Hauts-de-France region, Inria, the MEL, the I-Site ULNE regarding project RPILOTE-19-004-APPRENF, the French National Research Agency under PEPR IA FOUNDRY project (ANR-23-PEIA-0003). S.Vashishtha is partially supported by the ANR program “AI PhD@Lille”. O. Maillard acknowledges the Inria-Kyoto University Associate Team “RELIANT” for supporting the project. The authors thank Emilie Kaufmann for useful comments on the manuscript. S. Vashishtha thanks Ismaël Castillo for their lectures on Bayesian nonparametric statistics at the 51st Saint Flour Probability summer school, July 2023.
@@
Appendix A General Bayesian framework
@@
A general Bayesian modeling problem can be formulated as follows. We choose prior $Q$ on parameter $\Theta\in\mathbf{T}$ and the observation model $M$ as $P_{\Theta}$, observation space as $\mathbf{X}$. To summarize, both Bayesian and non-paramteric Bayesian models can be written as follows,
@@
For more clarity, the DP model can be re-written in the framework of Eqs. 12 as follows,
@@ A.2 Conjugacy
Appendix B Finite Stick breaking representation of Dirichlet Process priors
@@
We illustrate the application of Dirichlet processes for density estimation on a data set from the astronomy literature (Roeder, 1990). The measurements are velocities at which galaxies in the Corona-Borealis region are moving away from our galaxy. If the galaxies are clustered, the velocity density will be multimodal, with clusters corresponding to modes. This happens to be the case, and the multi-modal nature is evident in the CDF of the data in Figure 5 where the left and right regions of the CDF are almost flat, and most mass resides in the center. Starting with a $\mathrm{DP}(\alpha,N(0,1))$ prior, we obtain a DP posterior, and the spread of distributions sampled from the DP posterior (not shown) can be seen as confidence-set of the density estimate through Dirichlet process.
Figure 5: A random measure sampled from DP prior, DP posterior compared against original galaxy dataset distribution.
Appendix C DPPS for a Gaussian bandit
@@
Appendix D Choice of hyperparameters in numerical experiments
@@
Appendix E Running costs of DPPS
Here we detail the computational costs associated to a single-arm in each round. Let $n$ denote the number of observations for the arm. The important consideration in quantifying the running cost of DPPS is to scrutinize the posterior update step,
@@ Appendix E Running costs of DPPS
Appendix F Further related work
To the best of our knowledge, Dirichlet Processes in the context of bandits were first used in  (Clayton & Berry, 1985) to study a version of the single-arm Gittin’s index problem, when the probability distribution of the arm is assumed to be DP distributed. Use of Bootstrapping for Thompson sampling seems to have appeared first in Eckles & Kaptein (2014), which was further improved and made more systematic in  (Osband & Van Roy, 2015) where the authors also showed equivalence of Bootstrap-Thompson sampling (for Bernoulli-bandits) and Thompson sampling with Beta/Bernoulli priors in an exact sense, and speculated this equivalence for a wide class of bandit-environments if a proper mechanism for generating artifical history (or prior information) could be identified. As shown in the current paper, DPPS provides a neat and principled mechanism for incorporating prior information (or gnerating artificial history), and generalizes this equivalence. Non-Parametric Thompson sampling (NPTS) and Multinomial Thompson Sampling (TS) were introduced in (Riou & Honda, 2020) without highlighting any concrete Bayesian connection of the former algorithm. NPTS was adapted for robustness in  (Baudry et al., 2021). Some discussions concerning Bayesian interpretation of NPTS using DPs appeared in Belomestny et al. (2023) who provided a refined analysis of Multinomial TS. Aligning towards non-Bayesian side, a sample mean based algorithm guaranteeing $O(\log N)$ instance-dependent regret appeared in Agrawal (1995), a sub-sampling based algorithm was reported in Baransi et al. (2014) and analyzed for a two-arm bandit setting; a nonparametric Bootstrap based algorithm was reported in Kveton et al. (2019), and regret bounds derived for a Bernoulli bandit environment.
Appendix G Technical derivations
@@ G.2 Proof of Lemma 5
Define the random variable $X(t)=R_{t,a}-\mathbb{E}_{t}\left[R_{t,a}\right]$. Then, for arbitrary $\lambda\in\mathbb{R}$, applying Fact 4 to $\lambda X$ yields
@@ G.3 Proof of Lemma 2
where (b) follows from the Cauchy–Schwarz inequality, (c) follows from Fact 5, and (a) follows from Eq.11and (d) from the standard definition of mutual-information. ∎

### v2

@@ 6.2 A generalization of the information theoretic analysis in Russo & Van Roy (2016)
It’s easy to notice in the preceding analysis (specifically Eq. 11) that in the analysis of Russo & Van Roy (2016) there’s explicitly no restriction on $\mathbb{P}_{t}(A^{\star}=a)$ for it to be derived using a Bayes-rule based posterior-distributions of arm-rewards, $\mathbb{P}_{t}(R_{t,a})$, as is done in parametric Thompson sampling. However, this choice is rather implicit, given the decision theoretic and information theoretic coherency of Bayesian framework (Wald, 1961; Zellner, 1988). Moreover, Bayesian-framework is not limited to Bayes-rule based derivation of posterior distributions. Another valid Bayesian approach (Orbanz, 2009; Ghosal & van der Vaart, 2017) for obtaining posteriors is leveraging the property of conjugacy as discussed in Sec 3. In particular, most nonparametric priors do not satisfy the necessary conditions for Bayes rule (See section 8.1), and one relies on their conjugacy property to derive the corresponding posteriors. An elegant way to characterize the class of valid Bayesian approaches is based on martingales of probability measures, and we refer the interested reader to papers concerning predictive-Bayes  (Holmes & Walker, 2023; Fong et al., 2023; Fortini & Petrone, 2025) for details.
Admissible probability matching algorithms
From the preceding discussion, we can conclude that all probability matching algorithms which derive $\mathbb{P}_{t}(R_{t,a})$ (and hence $\mathbb{P}_{t}(A^{*}=a)$) using a valid Bayesian approach are admissible in the information theoretic analysis of Russo & Van Roy (2016). Additionally, these admissible algorithms would enjoy same bounds on their information-ratio and (consequently) Bayesian regret as those for parametric Thompson sampling (i.e. Lemma 1 and Theorem 2), if they satisfy certain auxiliary conditions required from the analysis. For the case of $\sigma$-sub Gaussian reward noise discussed before, we list these next.
Auxiliary conditions
It is easy to see that we require the following auxiliary conditions for an admissible probability matching algorithm to enjoy same bounds on Bayesian regret as those for parametric Thompson sampling in (Russo & Van Roy, 2016): In each round $t$,
  1. 1.
    The instantaneous reward noise, $X(t)$, in Lemma 2, is a $\sigma$-sub-Gaussian random variable,
  2. 2.
    $\texttt{KL}(\mathbb{P}_{t}(R_{t,a}|A^{\star}=a)||\mathbb{P}_{t}(R_{t,a}))$ in Lemma 2 is well defined.
The second condition holds if $\mathbb{P}_{t}(A^{\star}=a)>0$ owing to a classical fact in conditional probability,
Fact 2.
@@
(Williams, 1991) For any random variable $Z$ and event $E\subset\Omega$, where $\Omega$ is the probability space, if $\mathbb{P}_{t}(E)=0$, then $\mathbb{P}_{t}(E|Z)=0$ almost surely. Conversely, for any $x\in\mathcal{X}$ with $\mathbb{P}_{t}(X=x)>0$, $\mathbb{P}_{t}(Y|X=x)$ is absolutely continuous with respect to $\mathbb{P}_{t}(Y)$.
Next, we show that DPPS is an admissible probability matching algorithm that satisfies the auxiliary conditions listed above.
6.3 Bayesian regret of DPPS
@@
That DPPS is an admissible probability matching algorithm is easy to see: DPPS utilizes a valid Bayesian approach, i.e. conjugacy of DP priors/posteriors, to derive $\mathbb{P}_{t}(A^{\star}=a)$; Also DPPS satisfies both the auxiliary conditions discussed in previous sub-section; For condition (2), clearly, $\mathbb{P}_{t}(A^{\star}=a)>0$ for DPPS, whenever the base measure, $F_{0}$, of the DP prior (and hence of the corresponding DP posterior), $\texttt{DP}(\alpha,F_{0})$, is non-null. For condition (1), the following property of the tail of random measures sampled from DP priors/posteriors ensures $\sigma$-sub-Gaussian nature of the instantaneous reward noise, $X(t)$, whenever the base measure, $F_{0}$, of the DP prior, $\texttt{DP}(\alpha,F_{0})$, is $\sigma$-sub-Gaussian,
Fact 3 (Doss & Sellke (1982)).
@@
Therefore, DPPS is an admissible probability matching algorithm that satisfies the necessary auxiliary conditions under specific constraints on the DP priors, and hence whenever those constraints on the DP priors are satisfied, DPPS enjoys the same upper-bounds on Bayesian regret as those for parametric Thompson sampling. More concretely,
@@
For the setting of $\sigma$-sub-Gaussian arm reward distributions, starting with a DP-prior, $\texttt{DP}(\alpha,F_{0})$, with $F_{0}$ as a $\sigma$ sub-Gaussian distribution, the Bayesian regret of DPPS satisfies,
@@ 7 Conclusions and Perspectives
Firstly, we point that DPs are not the only Bayesian nonparametric priors on the space of distribution functions, and further generalization of DPPS is possible. For example, other probability matching algorithms using Pitman-Yor (Pitman & Yor, 2006) processes and Pólya-Tree priors  (Castillo, 2017; 2024) can be useful generalizations of DPPS. Next, we consider DPPS as a generic design principle, based on Bayesian non-parametric statistics, that can be extended to the setting of Markov Decision Processes as well, especially in the model-free scenario. Particularly, it seems that Bootstraped-DQN (Osband et al., 2016) can be interpreted as a DP based algorithm in the noninformative limit of the DP posteriors (i.e. Bayesian Bootstrap), and it would be interesting to extend it to a full-fledged DP implementation to account for any prior information (Osband et al., 2018) in a principled manner similar to that shown in this paper. Dirichlet Processes may also prove to be useful in efficiently representing value distributions in distributional reinforcement learning (Bellemare et al., 2023). We leave these intriguing research questions and extensions for future work.
@@ Acknowledgments
This work has been supported by the French Ministry of Higher Education and Research, the Hauts-de-France region, Inria, the MEL, the I-Site ULNE regarding project RPILOTE-19-004-APPRENF, the French National Research Agency under PEPR IA FOUNDRY project (ANR-23-PEIA-0003). S.Vashishtha is partially supported by the ANR program “AI PhD@Lille”. We thank Emilie Kaufmann and other members of our group (SCOOL) for encouraging discussions. S. Vashishtha thanks Ismaël Castillo for their lectures on Bayesian nonparametric statistics at the 51st Saint Flour Probability summer school, July 2023.
@@
8 General Bayesian framework
@@
A general Bayesian modeling problem can be formulated as follows. We choose prior $Q$ on parameter $\Theta\in\mathbf{T}$ and the observation model $M$ as $P_{\Theta}$, observation space as $\mathbf{X}$. To summarize, both Bayesian and non-parametric Bayesian models can be written as follows,
@@
For more clarity, the DP model can be re-written in the framework of Eqs. 12-13 as follows,
@@ 8.2 Conjugacy
9 Finite Stick breaking representation of Dirichlet Process priors
@@
We illustrate the application of Dirichlet processes for density estimation on a data set from the astronomy literature (Roeder, 1990). The measurements are velocities at which galaxies in the Corona-Borealis region are moving away from our galaxy. If the galaxies are clustered, the velocity density will be multimodal, with clusters corresponding to modes. This happens to be the case, and the multi-modal nature is evident in the CDF of the data in Figure 5 where the left and right regions of the CDF are almost flat, and most mass resides in the center. We estimate the CDF of this density using DP priors. Starting with a $\mathrm{DP}(\alpha=2,\mathrm{F_{0}}=\mathrm{N}(10,1))$ DP prior, we obtain a DP posterior, and the spread of distributions sampled from the DP posterior can be seen as confidence-set of the CDF of density estimated through Dirichlet process.
Figure 5: DP prior and DP posterior compared against original galaxy dataset distribution.
10 DPPS for a Gaussian bandit with unknown means and variances
@@
11 Choice of hyperparameters in numerical experiments
@@
12 Computational costs of DPPS
Improved performance and flexibility of DPPS (and other Bootstrap based algorithms such as NPTS) does come with higher computational cost. For example, in the 6-arm Bernoulli bandit environments of horizon $T=10000$, average run-time (over 200 independent runs) of DPPS was around 18 seconds, whereas that of parametric TS (conjugate prior/posterior) was 2-3 seconds. For the 7 arm DSSAT bandit problem, DPPS takes around 20 seconds, NPTS takes around 16 seconds. Sec. 12 gives a detailed overview of the computational complexity of DPPS. All this said, this run time of DPPS can be significantly brought down by utilizing self-similarity (Ghosal, 2010) of DP posteriors and parallel computation of DP posteriors that a construction exploiting this self-similarity would enjoy, which we plan to do in future.
Next, we detail the computational costs associated to a single-arm in each round. Let $n$ denote the number of observations for the arm. The important consideration in quantifying the running cost of DPPS is to scrutinize the posterior update step,
@@ 12 Computational costs of DPPS
13 Further related work
To the best of our knowledge, Dirichlet Processes in the context of bandits were first used in  (Clayton & Berry, 1985) to study a version of the single-arm Gittin’s index problem, when the probability distribution of the arm is assumed to be DP distributed. Use of Bootstrapping for Thompson sampling seems to have appeared first in Eckles & Kaptein (2014), which was further improved and made more systematic in  (Osband & Van Roy, 2015) where the authors also showed equivalence of Bootstrap-Thompson sampling (for Bernoulli-bandits) and Thompson sampling with Beta/Bernoulli priors in an exact sense, and speculated this equivalence for a wide class of bandit-environments if a proper mechanism for generating artifical history (or prior information) could be identified. As shown in the current paper, DPPS provides a neat and principled mechanism for incorporating prior information (or generating artificial history), and generalizes this equivalence. Non-Parametric Thompson sampling (NPTS) and Multinomial Thompson Sampling (TS) were introduced in (Riou & Honda, 2020) without highlighting any concrete Bayesian connection of the former algorithm. NPTS was adapted for robustness in  (Baudry et al., 2021). Some discussions concerning Bayesian interpretation of NPTS using DPs appeared in Belomestny et al. (2023) who provided a refined analysis of Multinomial TS. In many inference problems, mixture and diffusion priors can be useful to approximate the performance of Bayesian nonparametric priors Müller & Mitra (2013), and a nice line of works performs Thompson sampling utilizing such (parametric) priors  (Hong et al., 2022; 2020; Kveton et al., 2024). Aligning towards non-Bayesian side, a sample mean based algorithm guaranteeing $O(\log N)$ instance-dependent regret appeared in (Agrawal, 1995), a sub-sampling based algorithm was reported in (Baransi et al., 2014) and analyzed for a two-arm bandit setting; a nonparametric Bootstrap based algorithm was reported in (Kveton et al., 2019), and regret bounds derived for a Bernoulli bandit environment. Finally, we would like to comment here that Information theoretic analysis presented in this paper can also be seen as a special case of well established PAC-Bayes approach, and we refer the reader to an excellent article (and references therein) for details  (Alquier et al., 2024).
14 Technical derivations
@@ 14.2 Proof of Lemma 2
Define the random variable $X(t)=R_{t,a}-\mathbb{E}_{t}\left[R_{t,a}\right]$. Then, for arbitrary $\lambda\in\mathbb{R}$, applying Fact 1 to $\lambda X$ yields
@@ 14.3 Proof of Lemma 1
where (b) follows from the Cauchy–Schwarz inequality, (c) follows from Fact 2, and (a) follows from Eq.11and (d) from the standard definition of mutual-information. ∎
15 Bandit Algorithm in Riou & Honda (2020)
Algorithm 2 Algorithm in Riou & Honda (2020)
1:Horizon $T\geq 1$, number of arms $K\geq 1$
2:for $k=1...K$do
3:     Set $R_{k}:=[1]$, and $N_{k}:=1$
4:end for
5:for $t=1...T$do
6:     for $k=1...K$do
8:     end for
10:     Pull arm $I(t)$ and observe reward $R_{t,I(t)}$.
12:end for